A transverse wave produced on a spring has a frequency of 200 Hz and travels along the length of a spring of 80 m in 0.5 seconds. The speed of the wave will be
1
Expert's answer
2020-04-23T12:52:55-0400
"v=\frac{s}{t}=\frac{80}{0.5}=160\frac{m}{s}"